OpenStudy (anonymous):

A candy bar box is in the shape of a triangular prism. The volume of the box is 2,400 cubic centimeters Part A: What is the height of the box? Show your work. (5 points) Part B: What is the approximate amount of cardboard used to make the sides of the candy box? Explain how you got your answer. (5 points)

OpenStudy (anonymous):

@ganeshie8 please help

OpenStudy (anonymous):

do you know the formula for a triangular prism ?

OpenStudy (anonymous):

for the volume of a triangular prism*

OpenStudy (anonymous):

v=1/2 a*c*h ?

OpenStudy (anonymous):

@Jdosio

OpenStudy (anonymous):

okay well just plug in the values that you know and solve for what you dont

OpenStudy (anonymous):

This might take me a little while hold on

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I tried but I cant figure it out

OpenStudy (anonymous):

@Jdosio

OpenStudy (anonymous):

im trying to figure it out myself give me a minute

OpenStudy (anonymous):

are you sure you are not given anymore information

OpenStudy (anonymous):

I have a picture

OpenStudy (anonymous):

okay i think you have to use a system of equations

OpenStudy (anonymous):

here

OpenStudy (anonymous):

-_- you didn't think to mention the large amount of missing variables you possessed...

OpenStudy (anonymous):

anyway just use the pythagorean theorem a^2 + b^2 = c^2

OpenStudy (anonymous):

a = 17 b= 16 an c is your height which you have to solve for